OpenStudy (dtan5457):

Help graphing y=-sin(x-3pi/2)+2

OpenStudy (dan815):

what does that x-3pi/2 part do?

OpenStudy (dan815):

do you know what y=sin(x) looks like

OpenStudy (dtan5457):

yeah. i did most of it, im just not sure about the phase shift, im trying to get a picture.

OpenStudy (dan815):

okay look at this

OpenStudy (dan815):

suppose you said, v=x-3pi/2

OpenStudy (dan815):

y=-sin(x-3pi/2)+2 y=-sin(v)+2, <---- lets graph this y vs v first

OpenStudy (dan815):